Please Help:
Find the payment necessary to amortize the loan.
$1600; 12% compounded quarterly; 8 quarterly payments
A) $227.99
B) $322.09
C) $227.93
D) $205.49

Answer by rfer(16322)

PMT=(L*i*(1+i)^n)/(1+i)^n-1
PMT=(1600*(.12/4)*(1+.12/4)^8)/(1+i)^8-1
PMT=(1600*.03*(1.03)^8)/(1.03)^8
PMT 48*1.26677/.26677
PMT=60.80/.26677=227.93 (C)